A small bar of an unknown metal was left on the bench. The density of the metal is 11.5 g/cm3. An X-ray diffraction experiment was conducted to measures the edge of the face-centered cubic unit cell as 4.06 x 10-10 m. Find the atomic weight of this metal and tentatively identify this metal.
1
Expert's answer
2020-09-16T05:11:44-0400

"Density=" "Mass \over Volume" "=11.5g\/cm^3"

"Volume(i^3)=" "(4.06\times 10^{-8}cm^3)^3"

"=6.692\times 10^{-23}cm^3"

"11.5g\/cm^3=" "Mass of unit cell \over 6.692\times 10^{-23}cm^3"

"\therefore" Mass of unit cell"=7.696\times 10^{-22}g"

Since there are "4" units per unit cell, the mass of one atom is mass of unit cell/ 4. Converting the mass of one atom to the mass of 1 mole therefore is;

"7.696\times 10^{-22}g\over 4atoms" "=" "1.924\times 10^{-22}g\over atom" "\times 6.022\times 10^{23} atoms"

"=115.86g\/mol"

The compound is thus "FeCO_3"